Two positive point charges, each 16 μC , lie along the x-axis at x = –0.15 m and x = +0.15 m . Find the electric field at the point (0, 0.22 m ) on the y-axis.
Answer #1

r1 = distance from charge at x = 0.15 = sqrt(0.152 + 0.222) = 0.266 m

r2 = distance from charge at x = - 0.15 = sqrt(0.152 + 0.222) = 0.266 m

\theta = tan-1(0.22/0.15) = 55.71

q = charge = 16 x 10-6 C

Electric field components along X-direction cancel out and net electric field along Y-direction is given as

Enet = k q Sin55.71/r12 + k qSin55.71/r22 = 2 (9 x 109) (16 x 10-6) Sin55.71/(0.266)2 = 3.4 x 106 N/C